Three moles of an ideal gas (CV = 5 cal deg-1mol-1) at 10.0 atm and 00 are
converted to 2.0 atm at 500. Find ∆E and ∆H for the change.
1
Expert's answer
2020-08-18T08:27:03-0400

1) "\Delta E=nCvdT"

n=3

T2=50oC=323,15K

T1=0oC=273.15K

"\Delta E=nCvdT=3\times5\times(323.15-273.15)=750 cal"


2)

"\Delta H=nCpdT=n(Cv+R)dT"

"\Delta H=n(Cv+R)dT=3\times(5+2)\times(323.15-273.15)=1050 cal"